# PETLP: A Privacy-by-Design Pipeline for Social Media Data in AI Research

### v1

@@ 3.2 Privacy-by-Design
Privacy-by-design – mandated in Article 25 GDPR and reinforced by accountability obligations in Article 24 – requires privacy safeguards form part of a project’s architecture from conception. PETLP operationalises this through DPIAs as default design tools spanning the entire research lifecycle. Researchers initiate DPIAs during pre-registration, update them when processing changes, and use them to guide decisions throughout every pipeline phase (Figure 1, Appendix E).
During pre-registration, researchers should consult comprehensive risk frameworks to identify potential privacy challenges early. With researchers likely to increasingly employ LLMs, Barberá (2025) provides particularly valuable guidance: privacy and data protection risks with recommended mitigations (pp. 28–42), risk identification methodologies (pp. 48–56), estimation and evaluation protocols (pp. 57–65), control and mitigation strategies (pp. 66–73), residual risk assessment (p. 74), and state-of-the-art tools and benchmarks as of March 2025 (pp. 97–101). Whilst LLM-focused, these frameworks offer transferable insights for broader AI applications. Consulting such resources during initial DPIA development ensures researchers anticipate risks systematically rather than discovering them retrospectively.
This approach creates significant efficiencies, particularly for researchers relying on legitimate interests under Article 6(1)(f). As the UK Information Commissioner’s Office clarifies, a comprehensive DPIA encompasses all elements of a Legitimate Interest Assessment (LIA) whilst providing deeper risk analysis (ICO 2023). The DPIA’s structured evaluation of purpose, necessity, and proportionality automatically satisfies the LIA’s three-part test (European Data Protection Board 2024a), eliminating redundant assessments whilst creating the documentary evidence required for ongoing compliance.
@@ 3.3 Extract
Recent jurisprudence offers encouraging signals for academic researchers. The Hamburg District Court’s decision in LAION v. Kneschke (2024) affirmed that creating AI training datasets (LAION-5B (Schuhmann et al. 2022)) through web scraping constitutes legitimate text and data mining (TDM) for scientific research purposes under the DSM Directive. Notably, the court held that platform terms cannot override statutory research exceptions for qualifying institutions – a principle codified in DSM Article 7(1).
However, this protection applies only to specific researchers under specific conditions. To navigate these variations, we develop a practice-informed typology of four extraction channels: platform-authorised access, user-mediated collection, third-party aggregation, and self-directed extraction. This typology, derived from regulatory analysis (what the law permits), platform architectures (what is technically feasible), and documented research practice (what researchers actually do), reveals that extraction legality depends on three intersecting factors: the researcher’s institutional status (qualifying research organisation or commercial entity), the research purpose (scientific or commercial), and the specific legal basis invoked (DSM Article 3, Article 4, or legitimate interest) (Appendix D, Figure 6).
3.3.1 PLATFORM-AUTHORISED ACCESS
@@
3.3.2 USER-MEDIATED DATA COLLECTION
@@
3.3.3 THIRD-PARTY AGGREGATION SERVICES
@@
This suggests that datasets on platforms like Academic Torrents, despite their widespread academic use, exemplify these converging legal risks. Following Reddit’s enforcement action against Pushshift, vast archives of Reddit data now circulate via torrents, typically without licenses beyond warnings that content ‘may be protected by copyright’ (Academic Torrent 2025). This situation presents researchers with an ethical paradox: datasets fundamental to research exist in legal limbo, simultaneously violating platform terms, database rights, and data protection law. That these datasets migrated from Pushshift to torrents following platform enforcement underscores their legally dubious status.
3.3.4 SELF-DIRECTED EXTRACTION
@@
For qualifying research organisations, DSM Article 3 provides robust protection. The LAION v. Kneschke court explicitly held that extracting publicly available image URLs for AI dataset creation constitutes legitimate TDM for scientific research. Platforms may implement measures to ensure the security and integrity of the networks and databases, but these ‘shall not go beyond what is necessary’ (Article 3(3)). Rate limiting might be proportionate; blanket automated access bans likely exceed necessity.
Commercial researchers face a sharply different landscape. Without Article 3 protection, they must rely on Article 4’s general TDM exception, which platforms can exclude through machine-readable opt-outs. Reddit’s robots.txt and ToS constitute such opt-outs, rendering commercial scraping legally impermissible.
@@ 3.4 Transform
3.4.1 DATA PREPROCESSING AND IP RIGHTS
Data transformation necessarily creates multiple reproductions falling within InfoSoc Article 2’s broad scope. Every preprocessing step – loading data into memory for cleaning, creating intermediate files during normalisation – constitutes reproduction requiring authorisation. Whilst Article 5(1) InfoSoc exempts ‘transient or incidental’ copies forming an integral and essential part of a technological process, transformation rarely qualifies. Cleaned datasets and extracted features have independent value beyond mere technical facilitation.
@@ 3.4.1 DATA PREPROCESSING AND IP RIGHTS
3.4.2 PRIVACY-PRESERVING TRANSFORMATION
@@
Article 5(1)(c) requires data minimisation: retaining only data ‘adequate, relevant and limited to what is necessary’. For transformation, this means actively removing unnecessary fields, aggregating where individual-level detail is not required, and resisting the temptation to preserve “potentially useful” data without clear purpose. The EDPB specifically expects ‘data minimisation strategies and techniques employed to restrict the volume of personal data included in the training process’ (European Data Protection Board 2024b, Paragraph 51), applicable even at preprocessing stages.
@@ 3.5 Load
The Load phase migrates transformed social media data into secure storage systems – databases, data warehouses, or cloud repositories – where it becomes accessible for querying and analysis. Unlike the Presentation phase where model training occurs, Load focuses exclusively on establishing compliant storage infrastructure and internal governance. Here, the safeguards defined in the Transform stage are encoded into concrete storage architectures, access protocols, and retention schedules (Appendix D, Figure 8).
3.5.1 DATA ACCESSIBILITY AND TRANSFER
Transformed data requires responsible access controls, consistent with the GDPR principles. Under the GDPR, fully anonymised data (Recital 26) falls outside its scope. However, the EDPB warns that achieving true anonymisation with social media data proves difficult (Article 29 Data Protection Working Party 2017), meaning most stored datasets remain pseudonymised and subject to transfer restrictions under Articles 45–49 GDPR:
  1. 1.
    Adequacy decisions (Article 45): Transfers to countries deemed by the European Commission to offer ‘essentially equivalent’ protection [^3] https://commission.europa.eu/law/law-topic/data-protection/international-dimension-data-protection/adequacy-decisions_en require no additional safeguards.
  2. 2.
    Appropriate safeguards (Article 46): Transfers to non-adequate countries (e.g., the US, unless covered by the EU-US Data Privacy Framework [^4] https://www.dataprivacyframework.gov/list) require Standard Contractual Clauses (SCCs), Binding Corporate Rules (BCRs), or administrative arrangements.
  3. 3.
    Derogations (Article 49): In exceptional circumstances, explicit consent or public interest justifications may be invoked – but only for non-repetitive, non-systematic transfers.
@@ 3.5.1 DATA ACCESSIBILITY AND TRANSFER
3.5.2 STORAGE ARCHITECTURE AND SECURITY IMPLEMENTATION
Security obligations arise immediately upon data storage. Loading data into persistent storage triggers Article 32’s mandate for security measures ‘appropriate to the risk’. The EDPB emphasises that storage systems must implement state-of-the-art protections regardless of complexity or cost (European Data Protection Board 2024c, Paragraph 7). Essential measures include, but not limited to, role-based access controls limiting data to authorised researchers and encryption both at rest (e.g., Advanced Encryption Standard AES-256) and in transit (e.g., Transport Layer Security TLS 1.3).
Cloud storage requires additional contractual safeguards. Major cloud providers offer pre-approved Data Processing Addenda meeting Article 28 requirements (Amazon Web Services, Inc. 2023; Google LLC 2025; Microsoft Corporation 2025). When loading data into these platforms, researchers should verify agreements cover their specific storage architecture.
3.5.3 RETENTION SCHEDULING AND DELETION PROTOCOLS
@@ 3.6 Present
3.6.1 DATA DISTRIBUTION
Re-identification Risks
Disseminating research findings requires careful alignment with the GDPR’s core principles – especially purpose limitation (Article 5(1)(b)) and data protection by design (Article 25). The Present stage focuses on how outputs may indirectly re-expose personal data, including through seemingly benign presentation methods.
@@ 3.6.1 DATA DISTRIBUTION
Transparency-Restriction Tensions
Public presentation requires navigating tensions between transparency mandates and platform restrictions. Reddit’s Developer Terms prohibit third-party data redistribution (Reddit, Inc. 2024c, Section 7.4), whilst major funders like Horizon Europe mandate open access to underlying data (European Innovation Council and SMEs Executive Agency 2024).
Whilst the LAION v. Kneschke decision affirmed that TDM copyright exceptions cover dataset creation, it did not extend this protection to distribution. The court’s narrow focus on reproduction rights leaves researchers without clear legal grounds to redistribute extracted content, even when initial extraction was lawful under DSM Article 3.
@@
3.6.2 AI MODEL DISTRIBUTION
Embedded Privacy Risks
@@ 3.6.2 AI MODEL DISTRIBUTION
The EDPB warns that ‘AI models trained on personal data cannot, in all cases, be considered anonymous’ (European Data Protection Board 2024b, Paragraph 34), meaning that model publication may constitute ongoing personal data processing. This concern is reinforced by the observation that ‘information from the training dataset, including personal data, may still remain absorbed in the parameters of the model’ (European Data Protection Board 2024b, Paragraph 31). Together, these warnings highlight that model release carries inherent privacy risks that persist beyond the training phase.
@@
Whilst differential privacy techniques such as Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al. 2016), implemented through tools like Opacus (Aketi et al. 2025), offer theoretical protection against membership inference, their practical deployment varies significantly by model scale. For smaller models (<100M parameters) typical of classification tasks, Opacus provides effective privacy protection with minimal utility loss. Medium-scale models like BERT (100M-1B parameters) maintain acceptable performance, particularly when fine-tuning rather than training from scratch. Recent advances including FSDP support, Ghost Clipping for memory efficiency, and LoRA for parameter-efficient fine-tuning now enable Opacus to scale to LLMs like Llama-3-8b. However, while the paper successfully demonstrates that these large models can be trained with differential privacy, Aketi et al. (2025) focus solely on demonstrating feasibility and computational efficiency, leaving the critical question of model quality and utility under privacy constraints unexplored for these large-scale deployments.
Copyright Liability
@@
Model publication introduces copyright risks beyond privacy concerns. We re-emphasise that the LAION v. Kneschke decision distinguished dataset creation from model training, explicitly declining to rule whether training and distributing AI models falls within TDM exceptions. This judicial restraint leaves researchers uncertain about model publication rights, even when dataset creation was lawful. Until the pending CJEU case (Like Company v. Google Ireland, C-250/25) provides clarity, researchers face three liability scenarios (Rosati 2024): (1) direct liability when published models generate outputs substantially reproducing training content; (2) secondary liability when third parties use published models to create infringing content; and (3) distribution liability for sharing datasets containing copyrighted material.
Importantly, contractual disclaimers accompanying research outputs provide limited protection. Whilst researchers might attempt to exclude their liability through such provisions (e.g., standardised open-source licences to specify liability limitations), these contractual arrangements typically cannot override statutory copyright protections (Rosati 2024). Legal liability may persist despite explicit disclaimers.
Contractual Constraints
@@
For social media research, platform-specific restrictions compound these risks. Reddit prohibits using its data for AI training without explicit permission (Reddit, Inc. 2024c, Section 4.2). Whilst qualifying research organisations may invoke DSM Article 3 for dataset creation, the uncertainty around model training and publication means researchers should implement technical measures to prevent models from reproducing training data verbatim, or consider publishing model architectures and training procedures rather than trained weights.
@@ 4 Discussion
Whilst PETLP successfully integrates complex legal requirements into a unified technical framework, practical limitations and emerging research directions warrant further development.
Implementation and Documentation Challenges: The framework lacks empirical validation – PETLP’s overheads from iterative DPIAs to differential privacy remain unquantified, limiting practical privacy-utility guidance. This gap reflects broader responsible AI governance challenges. Stakeholders struggle to operationalise transparency artifacts effectively; interviews with government and civil society actors reveal existing tools (Model Cards, Data Cards) prioritise machine-readability over human accessibility, failing GDPR Article 12(1)’s ‘plain language’ requirement (Kawakami, Wilkinson, and Chouldechova 2024). Future implementations must systematically address data provenance – documenting origin, transformations, and validity throughout the pipeline (Longpre et al. 2024) – whilst ensuring outputs remain actionable for both implementers and affected communities. This requires tools that contextualise privacy risks to specific use cases, bridging compliance documentation and practical utility.
Platform and Jurisdictional Complexity: Beyond implementation challenges, PETLP faces generalisation limitations. Our Reddit focus cannot capture platform-specific architectures and evolving policies. The European emphasis compounds complexity for global collaborations navigating what Arnold et al. (2024) term a ‘dizzying array’ of conflicting governance instruments. Extending PETLP demands adaptive mechanisms mapping these regulatory variations whilst maintaining core privacy principles.
Future Directions: Three developments could transform PETLP from conceptual framework to operational toolkit. First, automated provenance tools tracking data lineage whilst satisfying transparency requirements could qualify for regulatory safe harbours (Longpre et al. 2024). Second, jurisdiction-aware compliance modules leveraging AGORA could dynamically identify applicable regulations (Arnold et al. 2024). Third, privacy-utility benchmarks quantifying trade-offs between differential privacy and model performance would address the gap between theoretical protections and practical deployment – particularly imperative given evidence of personally identifiable information leakage in foundation models (Nasr et al. 2023). These enhancements would enable researchers to navigate evolving social media AI research whilst maintaining both compliance and scientific rigour.

### v2

@@ 3.2 Privacy-by-Design
Privacy-by-design – mandated in Article 25 GDPR and reinforced by accountability obligations in Article 24 – requires privacy safeguards form part of a project’s architecture from conception. PETLP operationalises this through DPIAs as default design tools spanning the entire research lifecycle. Researchers initiate DPIAs during pre-registration, update them when processing changes, and use them to guide decisions throughout every pipeline phase (Figure 1; see Appendix E for a comprehensive risk assessment example).
During pre-registration, researchers should consult comprehensive risk frameworks to identify potential privacy challenges early. Given the increasing adoption of LLMs in research contexts, Barberá (2025) offers particularly valuable guidance tailored to LLM privacy considerations: privacy and data protection risks with recommended mitigations (pp. 28–42), risk identification methodologies (pp. 48–56), estimation and evaluation protocols (pp. 57–65), control and mitigation strategies (pp. 66–73), residual risk assessment (p. 74), and state-of-the-art tools and benchmarks (as of March 2025) (pp. 97–101). Though LLM-focused, these frameworks provide transferable insights for broader AI applications, enabling researchers to anticipate risks during initial DPIA development rather than discovering them retrospectively.
This approach creates significant efficiencies, particularly for researchers relying on legitimate interests under Article 6(1)(f). As the UK Information Commissioner’s Office clarifies, a comprehensive DPIA encompasses all elements of a Legitimate Interest Assessment (LIA) whilst providing deeper risk analysis (ICO 2023). The DPIA’s structured evaluation of purpose, necessity, and proportionality automatically satisfies the LIA’s three-part test (European Data Protection Board 2024a), eliminating redundant assessments while creating the documentary evidence required for ongoing compliance.
@@ 3.3 Extract
Recent jurisprudence offers encouraging signals for academic researchers. The Hamburg District Court’s decision in LAION v. Kneschke (2024) affirmed that creating LAION-5B dataset (Schuhmann et al. 2022) through web scraping constitutes legitimate Text and Data Mining (TDM) for scientific research purposes under the DSM Directive. Notably, the court held that platform terms cannot override statutory research exceptions for qualifying institutions – a principle codified in DSM Article 7(1).
However, this protection applies only to specific researchers under specific conditions. To navigate these variations, we develop a practice-informed typology of four extraction channels: platform-authorised access, user-mediated collection, third-party aggregation, and self-directed extraction. This typology, derived from regulatory analysis (what the law permits), platform architectures (what is technically feasible), and documented research practice (what researchers actually do), reveals that extraction legality depends on three intersecting factors: the researcher’s institutional status (qualifying research organisation or commercial entity), the research purpose (scientific or commercial), and the specific legal basis invoked (DSM Article 3 and 4, and GDPR legal bases) (Appendix D, Figure 6).
Platform-authorised Access
@@
User-mediated Data Collection
@@
Third-party Aggregation Services
@@
This suggests that datasets on platforms like Academic Torrents, despite their widespread academic use, exemplify these converging legal risks. Following Reddit’s enforcement action against Pushshift, vast archives of Reddit data now circulate via torrents, typically without licenses beyond warnings that content ‘may be protected by copyright’ (Academic Torrent 2025). This situation presents researchers with an ethical paradox: datasets fundamental to research exist in legal limbo, potentially violating platform terms, database rights, and data protection law. That these datasets migrated from Pushshift to torrents following platform enforcement underscores their legally dubious status.
Self-directed Extraction
@@
For qualifying research organisations, DSM Article 3 provides robust protection. The LAION v. Kneschke court explicitly held that extracting publicly available image URLs for AI dataset constitutes legitimate TDM for research. Platforms may implement measures to ensure the security and integrity of the networks and databases, but these ‘shall not go beyond what is necessary’ (Article 3(3)). Rate limiting might be proportionate; blanket automated access bans likely exceed necessity.
Commercial researchers face a sharply different landscape. Without Article 3 protection, they must rely on Article 4’s general TDM exception, which platforms can exclude through machine-readable opt-outs. Reddit’s robots.txt and ToS constitute such opt-outs, rendering commercial scraping legally impermissible without prior approval.
@@ 3.4 Transform
Data Preprocessing and IP rights
Data transformation necessarily creates multiple reproductions falling within InfoSoc Article 2’s broad scope. Every preprocessing step – loading data into memory for cleaning, creating intermediate files during normalisation – constitutes reproduction requiring authorisation. While Article 5(1) InfoSoc exempts ‘transient or incidental’ copies forming an integral and essential part of a technological process, cleaned datasets and extracted features have independent value beyond mere technical facilitation.
@@
Privacy-Preserving Transformation
@@
Article 5(1)(c) requires data minimisation: retaining only data ‘adequate, relevant and limited to what is necessary’. For transformation, this means actively removing unnecessary fields, aggregating where individual-level detail is not required, and resisting the temptation to preserve potentially useful data without clear purpose.
@@ 3.5 Load
The Load phase migrates transformed social media data into secure storage systems – databases or cloud repositories – where it becomes accessible for querying and analysis. Unlike the Presentation phase where model training occurs, Load focuses exclusively on establishing compliant storage infrastructure and internal governance. Here, the safeguards defined in the Transform stage are encoded into concrete storage architectures, access protocols, and retention schedules (Appendix D, Figure 8).
Data Accessibility and Transfer
Transformed data requires responsible access controls, consistent with the GDPR principles. Under the GDPR, fully anonymised data (Recital 26) falls outside its scope. However, the EDPB warns that achieving true anonymisation with social media data proves difficult (Article 29 Data Protection Working Party 2017), meaning most stored datasets remain pseudonymised and subject to transfer restrictions under Articles 45–49 GDPR: (1) Adequacy decisions (Article 45) allow transfers to countries deemed by the European Commission to offer ‘essentially equivalent’ protection to require no additional safeguards; (2) Appropriate safeguards (Article 46) apply when transfers to non-adequate countries (e.g., the US, unless covered by the EU-US Data Privacy Framework) require Standard Contractual Clauses (SCCs), Binding Corporate Rules (BCRs), or administrative arrangements; and (3) Derogations (Article 49) permit, in exceptional circumstances, explicit consent or public interest justifications to be invoked – but only for non-repetitive, non-systematic transfers.
@@
Storage Architecture and Security Implementation
Security obligations arise immediately upon data storage. Loading data into persistent storage triggers Article 32’s mandate for security measures ‘appropriate to the risk’. The EDPB emphasises that storage systems must implement state-of-the-art protections regardless of complexity or cost (European Data Protection Board 2024c, Paragraph 7). Essential measures include, but not limited to, role-based access controls limiting data to authorised researchers, encryption both at rest (e.g., Advanced Encryption Standard AES-256) and in transit (e.g., Transport Layer Security TLS 1.3).
Cloud storage requires additional contractual safeguards. Major cloud providers offer pre-approved Data Processing Addenda meeting Article 28 requirements (Amazon Web Services, Inc. 2023; Google LLC 2025; Microsoft Corporation 2025; Supabase, Inc. 2025). When loading data into these platforms, researchers should verify agreements cover their specific storage architecture.
Retention Scheduling and Deletion Protocols
@@ 3.6 Present
Data Distribution
Disseminating research findings requires balancing scientific transparency against re-identification risks and platform redistribution restrictions.
Re-identification Risks Disseminating research findings requires careful alignment with the GDPR’s core principles – especially purpose limitation (Article 5(1)(b)) and data protection by design (Article 25). The Present stage focuses on how outputs may indirectly re-expose personal data, including through seemingly benign presentation methods.
@@
Transparency-Restriction Tensions Public presentation requires navigating tensions between transparency mandates and platform restrictions. Reddit’s Developer Terms prohibit third-party data redistribution (Reddit, Inc. 2024c, Section 7.4), whilst major funders like Horizon Europe mandate open access to underlying data (European Innovation Council and SMEs Executive Agency 2024).
Although the LAION v. Kneschke decision affirmed that TDM copyright exceptions cover dataset creation, it did not extend this protection to distribution. The court’s narrow focus on reproduction rights leaves researchers without clear legal grounds to redistribute extracted content, even when initial extraction was lawful under DSM Article 3.
@@
AI model Distribution
Publishing AI models introduces persistent risks as models may retain personal information, reproduce copyrighted content, and violate platform contractual obligations.
@@
Embedded Privacy Risks The EDPB warns that ‘AI models trained on personal data cannot, in all cases, be considered anonymous’ (European Data Protection Board 2024b, Paragraph 34), meaning that model publication may constitute ongoing personal data processing. This concern is reinforced by the observation that ‘information from the training dataset, including personal data, may still remain absorbed in the parameters of the model’ (European Data Protection Board 2024b, Paragraph 31). Together, these warnings highlight that model release carries inherent privacy risks that persist beyond the training phase.
@@
While differential privacy techniques such as Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al. 2016), implemented through tools like Opacus (Aketi et al. 2025), offer theoretical protection against membership inference, their practical deployment varies significantly by model scale. For smaller models (<100M parameters) typical of classification tasks, Opacus provides effective privacy protection with minimal utility loss. Medium-scale models like BERT (100M-1B parameters) maintain acceptable performance, particularly when fine-tuning rather than training from scratch. Recent advances including FSDP support, and LoRA for parameter-efficient fine-tuning now enable scaling to LLMs like Llama-3-8b. However, Aketi et al. (2025) focus solely on demonstrating feasibility and computational efficiency for large models, leaving the critical question of model quality and utility under privacy constraints unexplored.
@@
Copyright Liability Model publication introduces copyright risks beyond privacy concerns. We re-emphasise that the LAION v. Kneschke decision distinguished dataset creation from model training, explicitly declining to rule whether training and distributing AI models falls within TDM exceptions. This judicial restraint leaves researchers uncertain about model publication rights, even when dataset creation was lawful. Until the pending CJEU case (Like Company v. Google Ireland, C-250/25) provides clarity, researchers face three liability scenarios (Rosati 2024): (1) direct liability when published models generate outputs substantially reproducing training content; (2) secondary liability when third parties use published models to create infringing content; and (3) distribution liability for sharing datasets containing copyrighted material. Importantly, contractual disclaimers accompanying research outputs provide limited protection. While researchers might attempt to exclude their liability through such provisions (e.g., standardised open-source licences to specify liability limitations), these contractual arrangements typically cannot override statutory copyright protections (Rosati 2024). Legal liability may persist despite explicit disclaimers.
@@
Contractual Constraints For social media research, platform-specific restrictions compound these risks. Reddit prohibits using its data for AI training without explicit permission (Reddit, Inc. 2024c, Section 4.2). Although qualifying research organisations may invoke DSM Article 3 for dataset creation, the uncertainty around model training and publication means researchers should implement technical measures to prevent models from reproducing training data verbatim, or consider publishing model architectures and training procedures rather than trained weights.
@@ 4 Discussion
While PETLP successfully navigates overlapping legal regimes as validated through our Reddit case study, it cannot resolve the structural tensions inherent in social media AI research governance.
From Checkbox to Continuous Compliance: Traditional AI governance artifacts – Model Cards, Data Cards, Transparency Notes – often deliver only “checkbox compliance”, satisfying regulatory requirements while failing to enable meaningful oversight (Kawakami, Wilkinson, and Chouldechova 2024). PETLP’s living DPIA approach embeds compliance decisions directly into research workflows, shifting from performative to operational transparency. This treats privacy as a design principle shaping methodological choices from extraction through deployment. However, even this approach operates within an ecosystem where transparency often legitimises rather than scrutinises AI systems.
Navigating Fragmented Governance: The regulatory landscape lacks coherence, with governance instruments emerging daily from disparate actors (Arnold et al. 2024). PETLP structures this complexity through decision trees, yet cannot eliminate underlying fragmentation. The LAION-5B case illustrates a related challenge. Even when datasets satisfy one legal domain (DSM Article 3), they may violate others (the dataset’s deprecation following discovery of illegal content (Thiel, Stroebel, and Portnoff 2023; NeurIPS 2025)). Researchers must therefore navigate not just fragmented governance but also the gaps between legal domains – where copyright compliance, platform terms, content regulations, and data protection operate independently.
Platform Heterogeneity as Structural Barrier: While LAION v. Kneschke protects academic scraping legally, technical enforcement varies unpredictably. Reddit’s six-month API limitation constrains longitudinal research; X’s pricing excludes unfunded researchers; Meta’s account suspensions create chilling effects. PETLP acknowledges this through four extraction channels with distinct trade-offs, yet each platform demands bespoke implementation, multiplying complexity for multi-platform studies.
Operationalising Responsible Practice: PETLP requires complementary infrastructure for practical deployment. First, institutional support through template DPIAs and integrated legal-ethical review would reduce researcher burden. Universities could extend IRB processes to encompass DPIA assessment. Second, shared computational environments with pre-configured differential privacy would enable wider adoption of privacy protections, addressing gaps between theoretical safeguards and practical deployment given documented PII leakage (Nasr et al. 2023). While infrastructures like European Open Science Cloud (EOSC) demonstrate federated research platforms, they currently lack privacy tooling. Third, automated compliance tools could bridge legal complexity and research practice. Provenance tracking systems that generate GDPR-compliant documentation while qualifying for regulatory safe harbours (Longpre et al. 2024) would transform PETLP’s decision trees into executable compliance pipelines. A forthcoming RedditHarbor toolkit demonstrates one path forward – automating PETLP’s compliance logic specifically for Reddit research. These developments recognise that responsible AI research requires not just frameworks but ecosystems where institutional support, technical infrastructure, and automated tools converge to make compliance achievable rather than aspirational.